OpenStudy (anonymous):

Help Please! What is the solution of the system? Use substitution. y = x + 6 y = 2x

OpenStudy (anonymous):

By substitution, we obtain: 2x = x + 6 2x - x = x - x + 6 x = 6

OpenStudy (anonymous):

(6,12) this is what i got